Select the set of equations that represents the following situation: The cost of six CDs and twelve DVDs is $73.26; and the cost of thirteen DVDs and twelve CDs is $140.11. How much do a CD and a DVD cost?
6x + 12y = 73.26; 13x + 12y = 140.11
6x • 12y = 73.26; 12x • 13y = 140.11
6x • 12y = 73.26; 13x • 12y = 140.11
6x + 12y = 73.26; 12x + 13y = 140.11

Answer by stanbon(75887)

Select the set of equations that represents the following situation: The cost of six CDs and twelve DVDs is $73.26; and the cost of thirteen DVDs and twelve CDs is $140.11. How much do a CD and a DVD cost?
6x + 12y = 73.26; 13x + 12y = 140.11
6x • 12y = 73.26; 12x • 13y = 140.11
6x • 12y = 73.26; 13x • 12y = 140.11
6x + 12y = 73.26; 12x + 13y = 140.11
----------------------------
6x + 12y = 73.26
12x + 13y = 140.11
----------------------------
Multiply thru 1st by 2
12x + 24y = 2*73.26
12x + 13y = 140.11
-----------
Subtract and solve for "y":
9y = 6.41
y = 0.7122 dollars (price of one DVD)
----
Solve for "x":
6x + 12*0.7122 = 73.26
6x = 61.71
x = $10.78 (price of one CD)
============
Note: These prices don't make much sense.
==============
Cheers,
Stan H.